I'm confused with this question because the wind is involved.
Pls. help me!
A plane travels 1200 km distance in 12 hrs with the wind. The return trip against the wind takes 15 hrs. Find the rate of the plane in still air.

Answer by checkley77(12569)

D=RT OR R=D/T
R=1200/12
R=100 MPH. IS THE SPEED WITH THE WIND
R=1200/15
R=80 MPH IS THE SPEED AGAINST THE WIND.
(100+80)/2=180/2=90 MPH IS THE SPEED OF THE PLANE IN STILL AIR.